OpenStudy (anonymous):

Solve for x: 5x + 2 = 4x − 9 a) 7 b) −7 c) 11 d) −11

OpenStudy (anonymous):

its going to be -11 because u first take the 4x and subtract it with the 5x to get x+2=-9 then u subtract the 2 to the -9 and u will get x=-11

OpenStudy (anonymous):

thanks for the medal

OpenStudy (anonymous):

youre welcome

OpenStudy (anonymous):

5x + 2 - 2 = 4x − 9 - 2 5x + 0 = 4x - 11 5x = 4x − 11 5x -4x = 4x -4x − 11 1x = 0-11 x=-11